OpenStudy (anonymous):

3 times the measure of an angle is 14 less than the measure of its complement. What is the measure of the angle? A.19° B.57° C.71° D.76°

OpenStudy (anonymous):

Do you know how to start?

OpenStudy (anonymous):

You there?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

Okay, So let x represent the angle.

OpenStudy (anonymous):

supplement = 180-x complement = 90-x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3x=(90-x)-14

OpenStudy (anonymous):

x=19

OpenStudy (anonymous):

ok. thanx

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

What did you do when you got 3x = 90 - x- 14

OpenStudy (anonymous):

For 3x = 90 - x- 14, x=19

OpenStudy (anonymous):

i mean like what did you do with the 14